# Towards Controllable Speech Synthesis in the Era of Large Language Models: A Survey

### v1

@@ IV-B Control Strategies
The control strategies in existing controllable TTS can be broadly classified into four categories: style tagging using discrete labels, speech reference prompt for customizing a new speaker’s voice with just a few seconds of voice input, controlling speech style using natural language descriptions, and the instruction-guided mode. We illustrate taxonomies of controllable TTS from the perspective of control strategies in Fig. 3.
@@
IV-B1 Style Tagging
This paradigm typically employs target control attributes, primarily emotion-related controls, as categorical label inputs to enable controllable speech synthesis. StyleTagging-TTS [223] utilizes a short phrase or word to represent the style of an utterance and learns the relationship between linguistic embedding and style embedding space by a pre-trained language model.
@@ IV-B1 Style Tagging
However, these methods are limited in expressive diversity, as they can only model a small set of pre-defined styles.
IV-B2 Reference Speech Prompt
@@
This paradigm aims to customize a new speaker’s voice with just a few seconds of voice prompt. The architecture can be abstracted into two main components: a speaker encoder that processes the reference speech and outputs a speaker embedding, and a conditional TTS decoder that takes both text and speaker embedding as input to generate speech that matches the style of the reference prompt. MetaStyleSpeech [225] and StyleTTS [88] use adaptive normalization as a style conditioning method, enabling robust zero-shot performance. GenerSpeech [90] introduces a multi-level style adapter to improve zero-shot style transfer for out-of-domain custom voices. SC VALL-E [205] facilitates control over synthesized speech’s emotions, speaking styles, and various acoustic features by incorporating style tokens and scale factors. ArtSpeech [236] revisits the sound production system by integrating articulatory representations into the TTS framework, improving the physical interpretability of articulation movements.
@@
To enhance the learning of contextual information and address the challenge of limited voice data from the target speaker, CCSP [237] proposes a contrastive context-speech pretraining framework that learns cross-modal representations combining both contextual text and speech expressions. DEX-TTS [183] separates styles into time-invariant and time-variant components, enabling the extraction of diverse styles from expressive reference speech. StyleTTS-ZS [239] leverages distilled time-varying style diffusion to capture diverse speaker identities and prosodies.
@@ IV-B3 Natural Language Descriptions
Recent studies explore controlling speech style using natural language descriptions that include attributes such as pitch, gender, and emotion, making the process more user-friendly and interpretable. In this paradigm, several speech datasets with natural language descriptions [101, 247, 106] and associated prompt generation pipelines [247, 102, 252] have been proposed. Detailed information about these datasets will be discussed in Section V. PromptTTS [101] uses manually annotated text prompts to describe five speech attributes, including gender, pitch, speaking speed, volume, and emotion. InstructTTS [105] introduces a three-stage training procedure to capture semantic information from natural language style prompts and adds further annotation to the NLSpeech dataset’s speech styles. PromptStyle [229] constructs a shared space for stylistic and semantic representations through a two-stage training process. TextrolSpeech [247] proposes an efficient prompt programming methodology and a multi-stage discrete style token-guided control framework, demonstrating strong in-context capabilities. NansyTTS [240] combines a TTS trained on the target language with a description control model trained on another language, which shares the same timbre and style representations to enable cross-lingual controllability.
@@
Considering not all details about voice variability can be described in the text prompt, PromptTTS++ [233] and PromptSpeaker [251] tries to construct text prompts with more details. PromptTTS 2 [102] designs a variation network to capture voice variability not conveyed by text prompts. ControlSpeech [106] proposes the Style Mixture Semantic Density (SMSD) module, incorporating a noise perturbation mechanism to tackle the many-to-many problem in style control and enhance style diversity.
@@
Other works also focus on improving controllability in additional aspects, such as the surrounding environment. Audiobox [187] introduces both description-based and example-based prompting, integrating speech and sound generation paradigms to independently control transcript, vocal, and other audio styles during speech generation. VoiceLDM [231] and AST-LDM [234] extend AudioLDM [253] to incorporate environmental context in TTS generation by adding a content prompt as a conditional input. Building on VoiceLDM, MS2KU-VTTS [242] further expands the dimensions of environmental perception, enhancing the generation of immersive spatial speech.
IV-B4 Instruction-Guided Control
@@
The natural language description-based TTS methods discussed above require splitting inputs into content and description prompts, which limits fine-grained control over speech and does not align with other AIGC models. VoxInstruct [103] proposes a new paradigm that extends traditional text-to-speech tasks into a general human instruction-to-speech task. Here, human instructions are freely written in natural language, encompassing both the spoken content and descriptive information about the speech. To enable automatic extraction of the synthesized speech content from raw text instructions, VoxInstruct uses speech semantic tokens as an intermediate representation, bridging the gap in current research by allowing the simultaneous use of both text description prompts and speech prompts for speech generation.
V Datasets and Evaluation
TABLE V: A summary of open-source datasets for controllable TTS.
@@ V Datasets and Evaluation
Dataset Hours #Speakers Labels Lang Release Time
Pit. Ene. Spe. Age Gen. Emo. Emp. Acc. Top. Des. Env. Dia.
Taskmaster-1 [254] / / en 2019.09
Libri-light [255] 60,000 9,722 en 2019.12
AISHELL-3 [256] 85 218 zh 2020.10
ESD [257] 29 10 en,zh 2021.05
GigaSpeech [258] 10,000 / en 2021.06
WenetSpeech [259] 10,000 / zh 2021.07
PromptSpeech [101] / / en 2022.11
DailyTalk [260] 20 2 en 2023.05
TextrolSpeech [247] 330 1,324 en 2023.08
VoiceLDM [231] / / en 2023.09
VccmDataset [106] 330 1,324 en 2024.06
MSceneSpeech [261] 13 13 zh 2024.07
SpeechCraft [262] 2,391 3,200 en,zh 2024.08
Abbreviations: Pit(ch), Ene(rgy)=volume=loudness, Spe(ed)=duration, Gen(der), Emo(tion), Emp(hasis), Acc(ent), Dia(logue), Env(ironment), Des(cription).
TABLE VI: Common objective and subjective evaluation metrics
@@
MCD [263] Objective Acoustic similarity
PESQ [264] Objective Perceptual quality
WER [265] Objective Intelligibility
MOS [266] Subjective Preference
CMOS [267] Subjective Preference
GT: Ground truth.
V-A Datasets
Achieving fully controllable TTS requires large-scale datasets with rich diversity and fine-grained annotations. In this subsection, we categorize speech datasets into four types according to the labels they provide, i.e., attribute tags such as age and gender, description, environment, and dialogues.
Tag-based speech datasets. Tag-based datasets [255, 256, 257, 258, 259, 260, 261] are specialized collections of speech data that include metadata tags, such as pitch, emotion, and energy, age, gender, and so on, to enhance the expressiveness and control of TTS systems. These datasets not only provide audio-text pairs but also offer additional labels that guide the model in generating diverse and context-aware speech outputs.
Description-based speech datasets. Description-based TTS datasets [101, 247, 231, 106, 262] contain speech data paired with detailed textual descriptions of the speech attributes or characteristics, such as emotion, speed, intonation, and style. These datasets enable training TTS models that interpret descriptive prompts to generate expressive and context-aware speech.
Speech environment datasets. Speech environment datasets [231, 268] are collections of speech data annotated with environmental labels, such as park, stadium, office, or street, capturing the acoustic characteristics of various real-world settings. These datasets are crucial for training models to handle diverse acoustic scenarios, improving their robustness and adaptability.
Dialogue speech datasets. Dialogue speech datasets [260, 254] capture conversational speech between two or more people, focusing on the natural flow, turn-taking, and context of human dialogue. These datasets are crucial for training systems in applications like chatbots, virtual assistants, and interactive storytelling.
We summarize open-source speech datasets for controllable TTS in Table V.
V-B Evaluation
@@
Objective evaluation metrics. Objective metrics offer automated and reproducible evaluations. Mel Cepstral Distortion (MCD) [263] measures the spectral distance between synthesized and reference speech, reflecting how closely the generated audio matches the target in terms of acoustic features. For intelligibility, the Word Error Rate (WER) [265] is used, comparing transcriptions of synthesized speech to the input text via automated speech recognition. Perceptual Evaluation of Speech Quality (PESQ) [264] is another objective metric designed to evaluate speech quality by comparing degraded audio with a clean reference. It is widely used in telecommunications and speech synthesis, PESQ models human auditory perception, producing a score (typically 1–4.5) that reflects intelligibility and distortion under various conditions, including noise or compression.
@@
Subjective evaluation metrics. The Mean Opinion Score (MOS) [266] is the most commonly used subjective metric. In MOS evaluations, listeners rate the naturalness of synthesized speech on a scale (e.g., 1 to 5), where higher scores indicate better quality. MOS captures human perception effectively but requires substantial resources for large-scale evaluations. Comparison Mean Opinion Score (CMOS) [267] further evaluates relative quality differences between two TTS audio samples. Participants listen to paired samples and rate their preference on a scale (e.g., -3 to +3, where negative values favor the first sample). CMOS is used to measure subtle improvements in TTS systems, complementing absolute MOS ratings.
Table VI summarizes common evaluation metrics for TTS.

### v2

@@ IV-C Control Strategies
The control strategies in existing controllable TTS can be broadly classified into four categories: style tagging using discrete labels or continuous control signals, reference speech prompt for customizing a new speaker’s voice with just a few seconds of voice input, controlling speech style using natural language descriptions, and instruction-guided speech attributes control. We illustrate taxonomies of controllable TTS from the perspective of control strategies in Fig. 4.
TABLE V: A summary of control strategies for TTS.
Control Strategy Control Signal
Style Tagging Discrete labels, continuous values, latent variables, speech audio
Speech Reference Prompt Speech audio
Natural Language Descriptions User input text, speech audio
Instruction-Guided Control User input text, speech audio
@@
Style Tagging. This paradigm enables controllable speech synthesis by adjusting key attributes such as pitch, energy, speech rate, and emotion. These attributes can be controlled using either categorical labels or continuous values. In this context, “tagging” refers to the process of assigning a control signal to a specific speech attribute. Some approaches utilize discrete labels to control synthesized speech. For instance, StyleTagging-TTS[182] employs short phrases or words to represent utterance styles. It learns the relationship between linguistic embeddings and style embeddings using a pre-trained language model. Emo-DPO[31] enables precise emotion control (e.g., “angry,” “happy”) via Direct Preference Optimization (DPO) [276] with LLM. By training on paired emotion-text data, it leverages contrastive learning to distinguish subtle prosodic differences between emotions. Users can specify an emotion label to shape the expressive quality of the generated speech. Spark-TTS [267] offers both coarse-grained control (e.g., gender, speaking style) and fine-grained adjustments (e.g., precise pitch values, speaking rate). Users can modify specially designed tokens through prompts to achieve customized speech synthesis with optional reference speech as input.
Other methods enable control by adjusting continuous input signals. DiffStyleTTS[221] models prosody hierarchically, incorporating coarse-grained implicit style conditions (extracted via GST [19]) and fine-grained explicit features. Users can adjust guiding scale factors to control pitch, energy, duration, and prosodic styles. DrawSpeech[222] provides an intuitive way to manipulate pitch and energy. Users can sketch rough prosody contours (e.g., rising or falling patterns), which the system refines into detailed contours through a sketch-to-contour predictor. A diffusion model then generates expressive speech that aligns with the user’s sketched prosody, offering precise control over vocal emphasis and intonation.
In addition, some approaches regulate speech attributes by modifying latent features rather than direct input signals. Cauliflow [186] controls speech rate and pausing by conditioning its flow-based duration model on user-defined parameters ($r_{s}$ for speed, $r_{p}$ for pause frequency) that adjust deviations from dataset averages during synthesis. DurIAN-E [195] allows users to manipulate expressive styles. It employs variance predictors for prosodic attributes and incorporates Style-Adaptive Instance Normalization (SAIN) layers, which dynamically adjust mel-spectrogram statistics based on predefined style embeddings. DiTTo-TTS [205], a diffusion transformer-based TTS model, removes dependencies on domain-specific features such as phonemes and durations while maintaining high performance. It controls speech rate by modifying the latent length predicted by a length predictor.
@@
These methods show great potential in controlling speech attributes by adjusting input signals or latent variables. However, these methods are limited in expressive diversity, as they can only model a small set of pre-defined attributes.
@@
Reference Speech Prompt. This paradigm aims to customize a new speaker’s voice with just a few seconds of voice prompt. The architecture can be abstracted into two main components: a speaker encoder that processes the reference speech and outputs a speaker embedding, and a conditional TTS decoder that takes both text and speaker embedding as input to generate speech that matches the style of the reference prompt. MetaStyleSpeech [184] and StyleTTS [88] use adaptive normalization as a style conditioning method, enabling robust zero-shot performance. GenerSpeech [90] introduces a multilevel style adapter to improve zero-shot style transfer for out-of-domain custom voices. SC VALL-E [242] facilitates control over synthesized speech’s emotions, speaking styles, and various acoustic features by incorporating style tokens and scale factors. ArtSpeech [210] revisits the sound production system by integrating articulatory representations into the TTS framework, improving the physical interpretability of articulation movements.
@@
To enhance the learning of contextual information and address the challenge of limited voice data from the target speaker, CCSP [211] proposes a contrastive context-speech pretraining (CCSP) framework that learns cross-modal representations, combining both contextual text and speech expressions. DEX-TTS [209] separates styles into time-invariant and time-variant components, enabling the extraction of diverse styles from expressive reference speech. StyleTTS-ZS [214] leverages distilled time-varying style diffusion to capture diverse speaker identities and prosodies.
@@
Natural Language Descriptions. Recent studies explore controlling speech style using natural language descriptions that include attributes such as pitch, gender, and emotion, making the process more user-friendly and interpretable. In this paradigm, several speech datasets with natural language descriptions [101, 243, 106] and associated prompt generation pipelines [243, 102, 273] have been proposed. Detailed information about these datasets will be discussed in Section V. PromptTTS [101] uses manually annotated text prompts to describe five speech attributes, including gender, pitch, speaking speed, volume, and emotion. InstructTTS [105] introduces a three-stage training procedure to capture semantic information from natural language style prompts and adds further annotation to the NLSpeech dataset’s speech styles. PromptStyle [190] constructs a shared space for stylistic and semantic representations through a two-stage training process. TextrolSpeech [243] proposes an efficient prompt programming methodology and a multi-stage discrete style token-guided control framework, demonstrating strong in-context capabilities. NansyTTS [215] combines a TTS trained on the target language with a description control model trained on another language, which shares the same timbre and style representations to enable cross-lingual controllability.
@@
Considering that not all details about voice variability can be described in the text prompt, PromptTTS++ [104] and PromptSpeaker [272] try to construct text prompts with more details. PromptTTS 2 [102] designs a variation network to capture voice variability not conveyed by text prompts. ControlSpeech [106] proposes the Style Mixture Semantic Density (SMSD) module, incorporating a noise perturbation mechanism to tackle the many-to-many problem in style control and enhance style diversity.
@@
Other works also focus on improving controllability in additional aspects, such as the surrounding environment. Audiobox [201] introduces both description-based and example-based prompting, integrating speech and sound generation paradigms to independently control transcript, vocal, and other audio styles during speech generation. VoiceLDM [194] and AST-LDM [203] extend AudioLDM [224] to incorporate environmental context in TTS by adding a content prompt as a conditional input. Building on VoiceLDM, MS2KU-VTTS [217] further expands the dimensions of environmental perception, enhancing the generation of immersive speech.
@@
Instruction-Guided Control. The description-based TTS methods discussed above require splitting inputs into content and description prompts, which limits fine-grained control over speech and does not align with other AIGC models. VoxInstruct [103] proposes a new paradigm that extends traditional text-to-speech tasks into a general human instruction-to-speech task. Here, human instructions are freely written in natural language, encompassing both the spoken content and descriptive information about the speech. To enable automatic extraction of the synthesized speech content from raw text instructions, VoxInstruct uses speech semantic tokens as an intermediate representation, bridging the gap in current research by allowing the simultaneous use of both text description prompts and speech prompts for speech generation. CosyVoice [17] introduces supervised semantic tokens derived from ASR models via vector quantization, enabling precise text-speech alignment. It employs an LLM for token generation and flow matching for synthesis. Controllability is enhanced through instruction fine-tuning, allowing adjustments in speaker identity, emotion, speaking rate, pitch, and paralinguistic elements (e.g., laughter, breath sounds) via textual instruction.
Some speech LLMs can also synthesize speech content, but offer less controllability. These speech LLMs are agents that take user instructions and call expert TTS models to synthesize speech content. AudioGPT [274] is a multimodal LLM designed to process and generate audio data. It extends the capabilities of general-purpose LLMs by incorporating audio/speech understanding, synthesis (FastSpeech2 [76]), and style conversion [90] modules. SpeechGPT [34] integrates speech and text via discrete representations. It employs a three-stage training strategy for cross-modal alignment. For speech synthesis control, it uses HiFi-GAN with speaker embeddings and chain-of-modality instructions, enabling instruction-guided speech generation, e.g., one can input “Please read the sentence: Today is a beautiful day.” FunAudioLLM [275] is an LLM enhancing voice interaction between humans and LLMs via SenseVoice [275] (multilingual speech recognition) and CosyVoice [17] (controllable speech synthesis). StepAudio [265] introduces a unified 130B speech-text model integrating understanding/generation, with a generative data engine enabling affordable voice cloning and a controllable TTS system (Step-Audio-TTS-3B [265]). Its speech synthesis achieves dynamic control through instruction-driven adjustments of dialects, emotions (anger/joy/sadness), singing styles, RAP vocals, and speaking rates.
V Datasets and Evaluation Metrics
TABLE VI: A summary of publicly available speech datasets for controllable TTS.
@@ V Datasets and Evaluation Metrics
Dataset Hours (at least) #Speakers (at least) Labels Lang Release Time
Pit. Ene. Spe. Age Gen. Emo. Emp. Acc. Top. Des. Dia.
IEMOCAP [277] 12 10 en 2008
RECOLA [278] 3.8 46 fr 2013
RAVDESS [279] / 24 en 2018
CMU-MOSEI [280] 65 1,000 en 2018
Taskmaster-1 [281] / / en 2019
AISHELL-3 [282] 85 218 zh 2020
Common Voice [283] 2,500 50,000 multi 2020
ESD [284] 29 10 en,zh 2021
GigaSpeech [285] 10,000 / en 2021
WenetSpeech [286] 10,000 / zh 2021
PromptSpeech [101] / / en 2022
MagicData-RAMC [287] 180 663 zh 2022
DailyTalk [288] 20 2 en 2023
TextrolSpeech [243] 330 1,324 en 2023
CLESC [289] $<$1 / en 2024
VccmDataset [106] 330 1,324 en 2024
MSceneSpeech [290] 13 13 zh 2024
Parler-TTS [273] 50,000 / en 2024
SpeechCraft [291] 2,391 3,200 en,zh 2024
Abbreviations: Pit(ch), Ene(rgy)=volume, Spe(ed), Gen(der), Emo(tion), Emp(hasis), Acc(ent), Top(ic), Des(cription), Dia(logue).
V-A Datasets
Fully controllable text-to-speech (TTS) systems require large-scale datasets that exhibit extensive diversity and include fine-grained annotations. Such datasets provide essential information that enables TTS models to generate highly expressive speech with precise control over various speech attributes.
In this subsection, we classify speech datasets into three categories based on the type of annotations they provide: (1) tag-based datasets, which include structured attribute labels such as age and gender; (2) description-based datasets, which contain detailed textual descriptions of speech characteristics; and (3) dialogue datasets, which capture natural conversational speech. Below, we provide an overview of these datasets and discuss their importance in advancing TTS research.
Tag-based Datasets. Tag-based datasets consist of speech recordings annotated with predefined attribute tags that characterize various aspects of the speech signal [282, 284, 285, 286, 288, 290, 287, 277, 278, 279, 280]. These attributes include pitch, energy, speed, age, gender, emotion, emphasis, accent, and topic, among others. Such datasets serve as a critical resource for developing TTS models with enhanced expressiveness and fine-grained control over speech generation. By incorporating attribute labels, models can be trained to adjust specific characteristics dynamically, enabling more personalized and context-aware speech synthesis.
Description-based Datasets. Description-based datasets go beyond structured labels by pairing speech samples with detailed textual descriptions that characterize various speech attributes, such as intonation, prosody, speaking style, and emotional nuances [101, 243, 106, 291, 273]. Unlike tag-based datasets, which provide predefined categorical labels, description-based datasets enable models to interpret nuanced, free-form textual prompts and generate speech that aligns with complex, context-dependent specifications. This category of datasets plays a crucial role in training TTS systems that can respond to natural language descriptions, making them highly suitable for applications requiring fine-grained expressiveness, such as storytelling, audiobooks, and personalized voice assistants.
Dialogue Datasets. Dialogue datasets [281, 288, 287] capture multi-turn conversational speech between two or more speakers, focusing on the natural flow of human interaction, including turn-taking, contextual dependencies, and prosodic variations. These datasets are particularly valuable for training conversational TTS models that can generate dynamic and contextually appropriate speech for chatbots and interactive applications. By incorporating dialogue-specific characteristics such as speaker intent, pauses, and conversational nuances, models trained on these datasets can produce more natural and engaging interactions for real-world scenarios.
By leveraging these different categories of speech datasets, researchers can develop more advanced and flexible TTS models capable of generating speech that is not only intelligible but also expressive, context-sensitive, and highly controllable. We summarize publicly available datasets in Table VI.
V-B Evaluation
TABLE VII: Common objective and subjective evaluation metrics.
@@ V-B Evaluation
MCD [292$\downarrow$ Objective Acoustic similarity
FDSD [293$\downarrow$ Objective Acoustic similarity
WER [294$\downarrow$ Objective Intelligibility
Cosine [295, 296$\downarrow$ Objective Speaker similarity
PESQ [297$\uparrow$ Objective Perceptual quality
SNR [298$\uparrow$ Objective Perceptual quality
MOS [299$\uparrow$ Subjective Preference
CMOS [300$\uparrow$ Subjective Preference
AB Test Subjective Preference
ABX Test Subjective Perceptual similarity
GT: Ground truth, $\downarrow$: Lower is better, $\uparrow$: Higher is better.
@@
Objective Evaluation Metrics. Objective metrics offer automated and reproducible evaluations. Mel Cepstral Distortion (MCD) [292] measures the spectral distance between synthesized and reference speech, reflecting how closely the generated audio matches the target in terms of acoustic features. A lower MCD value indicates a higher similarity between synthesized and reference speech, meaning better speech synthesis quality. Typically, an MCD value below 4 suggests good quality, while values above 6 may indicate significant distortion. The MCD is computed as follows:
$MCD=\frac{10}{\ln 10}\cdot\sqrt{2\sum_{d=1}^{D}(c_{d}^{(syn)}-c_{d}^{(ref)})^{2}},$ (3)
where $c_{d}^{(syn)}$ represents the d-th Mel Cepstral Coefficient (MCC) of the synthesized speech, $c_{d}^{(ref)}$ represents the d-th MCC of the reference speech, $D$ is the number of MCC, and $\frac{10}{\ln 10}\approx 4.342$ is a constant factor that converts the logarithm to a decibel scale.
Fréchet DeepSpeech Distance (FDSD) [293] is another metric designed to evaluate the quality and naturalness of synthesized speech. It is inspired by the Fréchet Inception Distance (FID) [301] used in image generation but adapted to speech by leveraging a deep speech recognition model. FDSD measures the statistical distance between the distributions of real (reference) and synthesized speech in the feature space of a pretrained speech recognition model, such as Deep Speech [302]. By comparing the mean and covariance of the extracted feature representations, FDSD provides a perceptually relevant assessment of speech synthesis quality. A lower FDSD means the synthesized speech is more similar to real speech. FDSD can be computed as:
$FDSD=||\mu_{s}-\mu_{r}||^{2}+\text{Tr}(\Sigma_{s}+\Sigma_{r}-2(\Sigma_{s}\Sigma_{r})^{1/2}),$ (4)
where $\mu_{s}$ and $\Sigma_{s}$ are the mean and covariance of the embeddings from the synthesized speech, $\mu_{r}$ and $\Sigma_{r}$ are the mean and covariance of the embeddings from the real (reference) speech, $||\mu_{s}-\mu_{r}||^{2}$ represents the squared Euclidean distance between the means, $\text{Tr}(\cdot)$ denotes the trace of a matrix, and $(\Sigma_{s}\Sigma_{r})^{1/2}$ is the geometric mean of the covariance matrices.
For intelligibility, the Word Error Rate (WER) [294] is used. It measures the difference between the recognized transcript and the reference transcript by computing the number of errors made in the transcription process. WER is computed as:
$WER=\frac{S+D+I}{N},$ (5)
where $S$ is the number of substitutions (wrong word in place of the correct word), $D$ is the number of deletions (missed words), $I$ is the number of insertions (extra words added), and $N$ is the total number of words in the reference transcript.
Cosine similarity (on speaker embeddings) measures similarity between speaker embeddings of synthesized and reference speech. It can be used to evaluate zero-shot TTS (voice cloning) methods, where higher values indicate better speaker similarity. Given two speaker embeddings, $\mathbf{e_{1}}$ and $\mathbf{e_{2}}$, their cosine similarity is defined as:
$CosSim(\mathbf{e_{1}},\mathbf{e_{2}})=\frac{\mathbf{e_{1}}\cdot\mathbf{e_{2}}}{\|\mathbf{e_{1}}\|\|\mathbf{e_{2}}\|},$ (6)
where speaker embeddings can be extracted from a pre-trained speaker embedding model (e.g., ECAPA-TDNN [295] and x-vectors [296]).
@@
Perceptual Evaluation of Speech Quality (PESQ) [297] is another objective metric designed to evaluate speech quality by comparing degraded audio with a clean reference. It is widely used in telecommunications and speech synthesis. PESQ models human auditory perception, producing a score in the range $[-0.5,-4.5]$ that reflects intelligibility and distortion under various conditions, including noise or compression. PESQ involves complex perceptual modeling, its core components can be summarized as:
$PESQ=a_{0}+a_{1}\cdot D_{frame}+a_{2}\cdot D_{time},$ (7)
where $D_{frame}$ is the frame-by-frame perceptual distortion, $D_{time}$ is the time-domain distortion, and $a_{0},a_{1},a_{2}$ are regression coefficients. One can refer to [297] for details.
Signal-to-Noise Ratio (SNR) measures the ratio of signal power to noise power. A higher SNR indicates a cleaner signal with less noise, while a lower SNR suggests that noise is dominating the signal. However, in TTS, noise can come from different sources, such as artifacts from vocoders, neural network distortions, or background noise in dataset recordings. A direct computation of SNR in TTS requires a reference clean speech signal ($x[n]$), a synthesized (or noisy) speech signal ($y[n]$), and extracting the noise component ($e[n]=y[n]-x[n]$) from the synthesized signal. The SNR for TTS systems can be computed as:
$SNR=10\log_{10}\left(\frac{P_{\text{signal}}}{P_{\text{noise}}}\right),$ (8)
where $P_{\text{signal}}=\frac{1}{N}\sum_{n=1}^{N}x[n]^{2}$ and $P_{\text{noise}}=\frac{1}{N}\sum_{n=1}^{N}e[n]^{2}$.
@@
Subjective Evaluation Metrics. The Mean Opinion Score (MOS) [299] is the most commonly used subjective metric. In MOS evaluations, listeners rate various aspects, such as naturalness, expressiveness, quality, intelligibility, et al., of synthesized speech on a scale from 1 to 5, where higher scores indicate better quality. MOS captures human perception effectively but is expensive for large-scale evaluations.
Comparison Mean Opinion Score (CMOS) [300] further evaluates relative quality differences between two TTS audio samples. Participants listen to paired samples and rate their preference on a scale (e.g., -3 to +3, where negative values favor the first sample). CMOS is used to measure subtle improvements in TTS systems, complementing absolute MOS ratings. MOS and CMOS scores are computed as the average scores across all listeners:
$MOS/CMOS=\frac{1}{N}\sum_{i=1}^{N}s_{i},$ (9)
where $s_{i}$ is the score given by the $i$-th listener, and $N$ is the number of listeners.
AB and ABX tests are also popular in evaluating TTS methods. An AB test involves presenting two versions of a synthesized speech (from different TTS models) to human listeners and asking them to choose which they prefer. The goal is to assess which model produces better-sounding speech based on certain criteria, such as naturalness, intelligibility, or clarity. In an ABX test, listeners compare two synthesized speech samples to a reference speech sample and determine which one is closer in terms of timbre, prosody, emotion, and other relevant features. ABX tests are widely used in evaluating zero-shot TTS methods. The AB/ABX test score for a model $m$ is:
$Score_{AB}/Score_{ABX}=\frac{N_{m}}{N},$ (10)
where $N_{m}$ represents the number of listeners who prefer the speech synthesized by model $m$, and $N$ denotes the total number of listeners.
Table VII summarizes widely used metrics for TTS.